OpenStudy (mackenzie2013):

In a recent vote between two nominees, 0.4 of the registered voters chose Candidate A and 1/4 of the registered voters chose Candidate B. What fraction of the total registered voters participated in the election?

OpenStudy (anonymous):

.4= 4/10 = 8/20 1/4 = 5/20 5/20 +8/20 =13/20